boohoo19:

A model rocket is launched with an initial upward velocity of 100 ft/s. The rocket's height h (in feet) after t seconds is given by the following. Find all values of for which the rocket's height is feet. h=100t-16t^2 Round your answer(s) to the nearest hundredth. (If there is more than one answer, use the "or" button.)

TurtlePrincess:

Hi what is the Rockers hight?

darkknight:

"Find all values of for which the rocket's height is feet. " you didn't copy paste correctly, we don't know what time to find for which height

Florisalreadytaken:

there is like a blank space i think -- thats why its blank

Florisalreadytaken:

oh nvm am wrong lol -- i kinda overthought it \[ h=100 \cdot \frac{25+\sqrt{625-4 h}}{8}-16 \cdot\left(\frac{25+\sqrt{625-4 h}}{8}\right)^{2} \] \[ \ \ \Updownarrow \] \[ h=h \] so there are infinite solutions -- sorry \( \underset{\huge\smile}{. .} \)

darkknight:

florisalreadytaken wrote:
oh nvm am wrong lol -- i kinda overthought it \[ h=100 \cdot \frac{25+\sqrt{625-4 h}}{8}-16 \cdot\left(\frac{25+\sqrt{625-4 h}}{8}\right)^{2} \] \[ \ \ \Updownarrow \] \[ h=h \] so there are infinite solutions -- sorry \( \underset{\huge\smile}{. .} \)
I think the question was originally asking to find all values of t when h = *insert height that they didn't give here*

Florisalreadytaken:

mhm i realised that : )